OpenStudy (anonymous):

Find the slope of the line by counting the change in y and the change in x. (Remember, slope is rise over run). (negative 3, 2), and (negative 2, negative 4) 2. Find the slope of the line by counting the change in y and the change in x. (Remember, slope is rise over run). (negative 3, 0) and ( 3, 2) 3. Find the slope of a line passing through the given points. (-4,2) and (1,-3) 4. Find the slope of a line passing through the given points. (3,3) and (4,0)

OpenStudy (anonymous):

(y2-y1)/(x2-x1) = the slope

OpenStudy (anonymous):

so (-4-2)/(-2-(-3))

OpenStudy (anonymous):

and so on with the other points, hope this helps :)

OpenStudy (anonymous):

.(2-0)/(3-(-3)) =2/6 =1/3